Titration of 20.00mh of a 1000m. • Find the pH during the thiethylamine (CH₃ CH2) 3N (K 5.2x10-4) 0.1000 M HCL
Answer #1

Since both solution have same molarity so equal volume of both solution will be required at equivalence point.

Et​​​​​3 ​​N ( aq) + HCl ( aq) \rightarrow Et​​​3​​NH +(aq) + Cl - ( aq)

Where Et = CH3CH2

Since Triethylamine is a weak base it's conjugate ion will hydolyse and make solution acidic

[ Et3NH + ( aq) ] = 20 mL x 0.1 M / ( 20 + 20 ) mL = 0.05 M

Et3NH​​ + ( aq )+ H2O ( l )\Leftrightarrow Et​​​​​ 3NH + ( aq )+ H3O + ( aq)

Hydrolsis constant ( Kh ) = Kw / Kb

= 10-14 / 5.2 x 10-4= 1.92 x 10-11

P​​​​​​H = (1/2 )( p​​​​​​kh - log c )

= ( 1/2 ) ( - log 1.92 x 10-11 - log 0.05 )

= 6.01